OpenStudy (anonymous):

Use the Pythagorean Theorem to solve for the hypotenuse in the right triangle: leg = 6, leg = 8, hypotenuse =

OpenStudy (trojanpoem):

 hypotenuse^2 = leg^2 + leg^2 hyp = sqrt( 8^2 + 8^2 = 8sqrt(2)

OpenStudy (anonymous):



OpenStudy (anonymous):

oh

OpenStudy (anonymous):

HAHHAHAHHAHHAHHAHHAHAHAHHAHAHAHAHHAHAHAHAHHAHAH

OpenStudy (anonymous):

CY BOOM

OpenStudy (mrnood):

so - how about using th epythagoras equation as required you know th e two smaller sides and you know (from pythaogoras) that H^2 = 6^2 +8^2 so work out H

OpenStudy (mrnood):

or maybe you could recognise the special case of a "3,4,5" triangle here?

OpenStudy (trojanpoem):

@terry1234566 , M4D ?